LOW-ENERGY PHOTOELECTRON IMAGING SPECTROSCOPY OF ANIONS

\underline{CHRISTOPHER L. ADAMS}, and J. MATHIAS WEBER, JILA, NIST, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309, USA; KENT M. ERVIN, Department of Chemistry, University of Nevada, Reno, 1664 N. Virginia St. MS 216, Reno, NV 89557-0216.

We report on photoelectron emission from nitroalkane anions using low-energy velocity map photoelectron imaging. The excess electron of these radical anions is weakly bound with adiabatic electronic affinities around 1500 cm^{−1} with relatively benign sensitivities to the length of the alkane chain. In addition to conventional photoelectron spectroscopy, we use vibrational excitation of the CH stretching vibration (2700-3000 cm^{−1}) to induce vibrational autodetachment. We have obtained an improved value for the adiabatic electron affinity (172± 6 meV), and definitively assigned the vibrational features in the photoelectron spectrum of the nitromethane anion. In addition, the vibrational autodetachment spectrum offers insight into vibrational modes involved in vibrational autodetachment.